Jefferson Lab Weekly Briefs May 9, 2012
Jefferson Lab Weekly Briefs
May 9, 2012
The latest "Montage" by Lab Director Hugh Montgomery has been posted to the website. In this newest article, titled "End of an Era," Mont reflects on the accomplishments of the lab during the 6 GeV era.
12 GeV Upgrade
An extended period of 12 GeV installation and commissioning, known as the Long Shutdown, is on schedule to commence on May 19. To highlight the importance of safety during this busy time, pre-shutdown meetings were held on May 7-8. These lab-wide meetings have become a staple of the kick-off process for scheduled machine downs. These meetings are beneficial as a tool to communicate the planned work activities and as a refresher course on the lab’s safety policies and procedures.
Hall A has been running g2p/GEp fairly smoothly, except for an 18-hour loss of time over the weekend due to a water pipe that broke in a power supply. The leak was fixed and several hoses were replaced, and the experiment was resumed.
The Q-weak experiment returned to production data taking with a 1.16 GeV beam. In addition to high-current parity data on the hydrogen target, the collaboration is making final calibration and consistency measurements with optics and background targets. In other, exciting news, Hall C Staff Scientist Silviu Covrig was named a winner of the 2012 Early Career Research Program award.
Scientists and Educators Team Up
Fellowship Helps Women Physicists Resume Paused Career
In Pictures: Doing Physics
The accelerator started out the week with quiet running, until instabilities in accelerator energy made delivering steady high-current beam to Hall C a challenge. Operations and Engineering support personnel began monitoring multiple signals to locate the source, but the randomness and the large number of possible causes worked against them. Several radiofrequency cavities were removed from service and even whole cryomodules of cavities were turned off, while previously idled modules were turned on in their place in an effort to isolate the source. As yet, a source has not been determined, but the search continues. On Wednesday, work was performed in the injector for the PEPPo experiment; Cryogenics replaced a failed cryogen tube that was iced over; and a waveguide and a klystron for two C100 superconducting cavities were replaced, among other maintenance items. Hall C was changed from 3-pass to 1-pass beam, and Hall A was changed from 1-pass to 2-pass beam. On Thursday, a problem with new software code resulted in one of the new C100 superconducting cryomodules being left off, with other modules turned back on to replace it, plus several hours of down time. On Friday, a spin dance was performed with Halls A and C, measuring the beam polarity (spin) at different injector settings to determine the optimal final setting. Because of the low-current running of Hall C, running was quiet over the weekend, with only minor changes to operating cavities or their gradients being necessary.
The FEL team completed the ultraviolet optics installation and began preparing the electron beam for UV operation. A number of systems were repaired and re-qualified, and a temperature bump on the quarter cryomodule was performed to drive out surface-absorbed helium.
Center for Theoretical and Computational Physics
A study of some fundamental aspects of the implementation of the so-called holographic description of QCD has been published by a collaboration between researchers in the Theory Center and the Instituto Balseiro in Argentina. This work addressed theoretical consistency requirements for such a description to be possible. The aim of the holographic picture of QCD is to describe QCD in its strong coupling regime, which is of key relevance for most of the hadronic phenomena being studied at JLab.
The alignment group continues to support the Hall A g2p experiment with numerous chicane moves. This work involves aligning the second FZ magnet, the harp assembly and the diagnostic girder to tolerances of less than 0.2 millimeters. Alignment also has quality inspected many of the magnets delivered for the 12 GeV Upgrade. Additionally, the group has begun to survey the network connecting Hall D and the Hall D tagger areas. Since there is no direct connection, or line of sight, between Hall D and where the tagger magnet is located, the survey network must be densified using the tagger area access ramp to the surface and then return to Hall D via its access ramp. The tolerance for this is in the quarter millimeter range. Cryomodule C100-7 was fiducialized, and the survey of the transfer line assembly for the Central Helium Liquefier-2 has been completed.
Environment, Safety, Health and Quality
Establishing an Electrically Safe Work Condition
Computing and Networking Infrastructure
Email Migrations into New System Ongoing
New Phone Number to Contact JLab Security
If the 911 call was dialed from a cell phone, you will also need to call JLab Security at 757-269-5822. This call is necessary to alert the security guards so they can be ready to escort emergency responders to the scene as quickly as possible.
Reminder: SAF100 Training Update for Everyone
JLab 2012 T-Shirts Arrive Next Week
Tornado Preparedness and Response Awareness Survey Results Posted
Tracking Thomas at JLab
Congratulations this week go to Pashupati Dhakal and Brita Hampton, who were the first to correctly identify the May 2 location. Honorable mentions go to Lyn Wells, Lori Zukerman, Samika Hawkins, Stephanie Vermeire, John Fischer, Mike Martin, Dan Moore, Ron Bartek, Andy Kowalski, Karen Bartek, Jim Follkie, Harry Fanning, Dick Owen, Subashini De Silva and Michele Solaroli. Check out the Tracking Thomas webpage for a better view of his last location and this week's new mystery photo.
JLab Calendar of Events
May 9-11: Science and Technology Review